### v1

# Advancing Brainwave-Based Biometrics: A Large-Scale, Multi-Session Evaluation

Abstract.
The field of brainwave-based biometrics has gained attention for its potential to revolutionize user authentication through hands-free interaction, resistance to shoulder surfing, continuous authentication, and revocability. However, current research often relies on single-session or limited-session datasets with fewer than 55 subjects, raising concerns about generalizability and robustness. To address this gap, we conducted a large-scale study using a public brainwave dataset of 345 subjects and over 6,000 sessions (averaging 17 per subject) recorded over five years with three headsets. Our results reveal that deep learning approaches outperform classic feature extraction methods by 16.4% in Equal Error Rates (EER) and comparing features using a simple cosine distance metric outperforms binary classifiers, which require extra negative samples for training. We also observe EER degrades over time (e.g., 7.7% after 1 day to 19.69% after a year). Therefore, it is necessary to reinforce the enrollment set after successful login attempts. Moreover, we demonstrate that fewer brainwave measurement sensors can be used, with an acceptable increase in EER, which is necessary for transitioning from medical-grade to affordable consumer-grade devices. Finally, we compared our findings with prior work on brainwave authentication and industrial biometric standards. While our performance is comparable or superior to prior work through the use of Supervised Contrastive Learning, standards remain unmet. However, we project that achieving industrial standards will be possible by training the feature extractor with at least 1,500 subjects. Moreover, we open-sourced our analysis code to promote further research.
Brainwave Authentication, EEG Authentication, Biometrics
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Biometricsccs: Security and privacy Usability in security and privacy
1. Introduction
Authentication is the cornerstone of securing digital interactions. While passwords have long been the dominant method, biometrics are increasingly preferred. According to the 2024 FIDO Survey [^1] https://fidoalliance.org/wp-content/uploads/2024/10/Barometer-Report-2024-Oct-29.pdf of 10,000 consumers, 28% preferred biometrics for signing into online accounts, apps, and smart devices, making it the most popular choice. Beyond established methods—such as fingerprint and face recognition— emerging biometric technologies are being explored to enhance usability and security. Brainwave authentication has gained attention for its hands-free operation, which enhances usability in devices like XR systems (Lin et al., 2018; Fallahi et al., 2024a). It resists shoulder surfing by relying on distinct neural activity patterns rather than observable actions or physical characteristics (Arias-Cabarcos et al., 2023). Additionally, brainwave-based systems can enable continuous authentication, improving security by monitoring user identity in active tracking (Nakanishi and Yoshikawa, 2015). Furthermore, research indicates that in cases of brainwave sample leakage, changing the stimulus can effectively reset the authentication pattern, thereby maintaining security (Lin et al., 2018). However, brainwave authentication is still in its early stages, and both permanence and performance remain significant challenges.
Prior studies on brainwave authentication mainly rely on limited public datasets or small, privately collected datasets. Consequently, studies either focus on single-session datasets (Schons et al., 2018; Bidgoly et al., 2022; Arias-Cabarcos et al., 2023; Fallahi et al., 2023) or small-scale multi-session datasets (Das et al., 2016; Wu et al., 2018; Maiorana, 2021; Debie et al., 2021). While single-session studies completely neglect session and temporal effects (Chaurasia et al., 2024; Huang et al., 2022), multi-session studies are limited to at most 55 (Debie et al., 2021) subjects and 270  (Maiorana and Campisi, 2017) sessions (6 per subject), increasing the risk of overfitting and restricting the ability to generalize results and identify key factors influencing performance and reliability over time. Moreover, the lack of publicly accessible code and data hinders reproducibility and further development, in contrast to the face recognition field, which benefits from open data and collaborative research.
To address this gap and better understand how different parameters affect the performance of brainwave authentication over time, we used the recently published PEERS dataset (Kahana et al., 2024), which comprises data from 345 subjects and over 6,000 sessions spanning five years. This dataset contains more than six times the number of subjects and 22 times the number of sessions compared to previous brainwave authentication evaluations. Our main contributions are as follows:
  • Benchmark (Sec 4.1): To identify the best-fit pipeline for multi-session authentication, we performed a comprehensive investigation of both classical and deep learning-based feature extraction approaches. Our results showed that pairwise-based learning methods, delivered superior performance for brainwave-based user authentication, reducing the EER from 27.3% to 10.9%. We also examined distance-based and binary classification-based feature comparison methods for brainwave data. While both approaches exhibited comparable performance, we chose the distance-based method because it does not require negative samples, making it more practical.
  • Analysis (Sec 4.2): To better understand the effective parameters influencing performance, we investigated the effects of test size, session intervals, and hardware type, as well. Our analysis reveals several key insights. First, a larger number of subjects in the test set does not necessarily increase the authentication error. However, smaller evaluation datasets introduce higher uncertainty due to the limited number of observations. Second, examining the time intervals between sessions shows that authentication errors increase over time: approximately 1% after two weeks, 4% after three months, and 12% after twelve months of enrollment. This indicates that, depending on the use case, it was beneficial to update the enrollment dataset after successful logins to reduce long-term error rates. Third, visualization of the embedding space reveals that samples from the same session are relatively close, which highlights the effect of sessions on the learning process, potentially caused by environmental noise, device setup, and changes in the subject over time. Finally, the analysis of the three devices used for dataset collection shows that cross-hardware authentication is only feasible when a sufficient number of subjects in the training set include cross-hardware sessions with the same devices; otherwise, the error would be too high.
  • Enhancing Performance (Sec 4.3): We investigated the trade-off between verification time and performance, reducing the error rate from 10.78% to 6.87% at the cost of increased verification time, which ranged from 1 to 32 seconds. Additionally, we examined the trade-off between enrollment sessions and error rates, observing that after the second enrollment session, the improvement in error rate diminished.
  • Channel Reduction (Sec 4.4): To evaluate the system’s performance under conditions resembling real-world scenarios, we reduced the number of channels in the PEERS dataset—collected using medical-grade devices—from 93 to 14, 7, and 4. These reductions correspond to sensor placements commonly found in consumer-grade EEG devices. The results showed a relatively small increase in error compared to the significant reduction in channels. For instance, the error increased from 10.78% EER with 93 channels to 13.97% with 14 channel. This modest increase can be attributed to the high correlation in EEG data, which helps mitigate the impact of channel reduction on performance.
  • Industrial Standard Comparison (Sec 4.5): We compare our results as the state of the art in brainwave authentication to industrial standards for biometric implementation, such as the National Institute of Standards and Technology (NIST) and the International Organization for Standardization (ISO). While our results outperform or are comparable to other brainwave authentication works, despite using a more challenging test set comprising 100 subjects, we observe that at the security levels expected in industrial standards, the FRR is high. However, our complementary experiments showed a linear relationship between the error rate and logarithm of the number of subjects in the feature extraction training set. Therefore, we predict that with at least 1,500 subjects, the standards could be met at a usable FRR.
  • Open Source: We published our source code, which, along with the public dataset, makes results easily reproducible and facilitates future error rate reductions by other researchers  [^2] https://github.com/kit-ps/NeuroShield/.
2. Background and Related Work
EEG measures electrical activity produced by neuronal firings in the human brain, typically recorded at the scalp using specialized sensors (Niedermeyer and da Silva, 2005; Hu and Zhang, 2019). German physiologist and psychiatrist Hans Berger recorded the first human EEG in 1924 for medical purposes, which remains one of its primary applications (Gupta et al., 2022; Adeli and Ghosh-Dastidar, 2010; Haas, 2003). Beyond medical use, EEG technology has advanced to facilitate direct communication between the brain and external devices, known as Brain-Computer Interfaces (BCI) (Kawala-Sterniuk et al., 2021). Recently, EEG data has also been found to include unique features linked to user identity, making it a promising method for user authentication (Gui et al., 2019).
Biometric authentication modalities should meet certain requirements to enable widespread usage. These requirements include universality (the modality must be present in all individuals), distinctiveness (it must uniquely identify an individual), permanence (it should remain consistent over time), and collectability (it must be measurable with available technology). Additional factors include performance (accuracy, speed, and resource efficiency), acceptability (user willingness to utilize the system), and circumvention resistance (difficulty in forging or bypassing) (Rui and Yan, 2018).
EEG is universal because each person’s brain produces EEG signals (Marcel and Millán, 2007). These signals exhibit distinctiveness due to individual brainwave patterns, making them uniquely identifiable (Kaliraman et al., 2019). Advances in technology are rapidly improving collectability, with new EEG headsets entering the market annually, offering better signal quality and usability (Yang and Deravi, 2017). However, without real-time implementation, it is difficult to comprehensively assess acceptability and circumvention resistance, though some research attempts have been made in this direction (Fallahi et al., 2024b; Wang et al., 2022).
Currently, the primary focus of research is on performance and permanence, as these are challenging aspects of EEG-based authentication. The two are closely interrelated, with performance metrics reflecting the system’s ability to maintain consistent accuracy over time. Factors such as hair growth, slight variations in sensor placement on the scalp, brain states (e.g., stress), and environmental conditions (e.g., noise) can significantly impact both performance and permanence in EEG-based authentication. However, most research studies evaluate the performance of brainwave authentication using single-session datasets (Schons et al., 2018; Bidgoly et al., 2022; Arias-Cabarcos et al., 2023; Fallahi et al., 2023), which neglect session invariance (permanence) completely. This limitation raises concerns that models may overfit to session-specific characteristics rather than capturing subject-unique features. Additionally, some studies (Debie et al., 2021; Chaurasia et al., 2024) consider short time intervals, such as within a few days, which may only partially account for the aforementioned issues. While a limited number of papers explore time intervals spanning more than a week, these studies (Das et al., 2016; Wu et al., 2018; Maiorana and Campisi, 2017; Maiorana, 2021) rely on non-public datasets and omit sharing their source code, hindering reproducibility and further research. Moreover, such studies have been conducted with a relatively small sample size and limited sessions (up to 54 subjects (Debie et al., 2021) and 270 sessions (Maiorana and Campisi, 2017)), which increases the risk of overfitting and restricts the generalizability of the results.
Therefore, a more comprehensive investigation of performance and permanence is needed using larger multi-session datasets to avoid overfitting. This would enable the examination of key parameters affecting performance over time and provide insights into improving brainwave authentication. Additionally, identifying the main challenges and ensuring reproducibility through open-source code sharing are essential for advancing the field.
3. Method
Biometric systems, including brainwave authentication, involve four key stages: data collection, preprocessing, feature extraction, and feature comparison, and, like other biometric systems, they operate in two phases: enrollment and verification. In the enrollment phase, brainwave signals are collected from the user to generate a reference profile, known as a template. During verification, performed after enrollment, brainwave signals are captured again and compared with the reference profile to confirm the user’s identity. Due to the high-dimensional and noisy nature of biometric data, including brainwaves, direct comparison of raw samples is impractical. Preprocessing is therefore applied post-data collection to eliminate noise and extract relevant features that preserve subject-specific uniqueness while reducing dimensionality. A similarity metric is subsequently employed to compare these processed features and determine whether to accept or reject the authentication request.
To address the challenges posed by the noisy and high-dimensional nature of brainwave data, some systems incorporate external stimuli during data collection (Alzahab et al., 2022; Arias-Cabarcos et al., 2023; Fallahi et al., 2023). These stimuli, such as visual or auditory signals, elicit distinct brainwave responses, allowing the system to timestamp and isolate signal segments directly related to the response. This approach enhances the signal-to-noise ratio and improves the reliability of the feature extraction process, ultimately strengthening the accuracy of the authentication system.
3.1. Dataset
Our study utilizes the Penn Electrophysiology of Encoding and Retrieval Study (PEERS) dataset (Kahana et al., 2024), provided by the University of Pennsylvania. Collected between 2010 and 2020, the dataset focuses on exploring EEG correlates of memory processes, particularly during tasks involving word memorization and recall. We used memorization tasks for user authentication, as the timestamps of word presentation available in the metadata can be considered as stimuli, while recall timestamps are not available. The dataset encompasses approximately 8.7 TB of EEG data, comprising recordings from 345 subjects across more than 6000 sessions. The participants range in age from 17 to 85 years, with an average age of 26.98 $\pm$ 15.86 years. The dataset includes 161 female and 136 male participants, while sex information for 48 subjects is unavailable. The PEERS dataset is particularly suited for our research as it includes multiple sessions per subject, enabling the investigation of inter-session variability, a critical factor for ensuring the robustness of real-world biometric systems. Data collection was conducted using three distinct medical grade headsets: the 129-channel Geodesic Sensor Net (GSN 200 model) [^2] https://med.stanford.edu/content/dam/sm/lucasmri/documents/16_0824_EGI_geodesic_sensor_net.pdf , the 129-channel HydroCel Geodesic Sensor Net. [^2] https://med.stanford.edu/content/dam/sm/lucasmri/documents/16_0824_EGI_geodesic_sensor_net.pdf, and the 128-channel BioSemi [^3] https://www.biosemi.com/headcap.htm headcap using the Biosemi ActiveTwo acquisition system. These diverse devices further enhance the dataset’s relevance by reflecting real-world variability in EEG acquisition systems.
3.2. Preprocessing
As the channel layout of the different headsets varied, we unified them by mapping the channels to a well-known extended version [^4] https://mne.tools/stable/auto_tutorials/intro/40_sensor_locations.html of the ten-twenty electrode system (Silverman, 1963), which includes 93 channels. Subsequently, the EEG data channels were reordered for each recording to align with the expected order during evaluation across all three devices. Following this, and in line with best practices for preprocessing in brainwave authentication (Chaurasia et al., 2024; Arias-Cabarcos et al., 2023; Peng, 2019; Maiorana, 2021), a common average reference is applied to the data, which helps in reducing noise and improving signal quality (Maiorana, 2021). Then, a bandpass filter is applied to retain frequencies between 0.1 Hz and 60 Hz. This filtering step removes low-frequency drift and high-frequency noise, preserving the components of the signal most relevant for EEG analysis (Peng, 2019; Felja et al., 2023). Finally, we extracted 1-second samples based on stimulus timestamps. The sample data was saved along with the session date and hardware type. In the paper, we refer to the oldest session for each subject as the enrollment sample, while the remaining sessions are referred to as verification samples, unless otherwise specified.
3.3. Feature Extraction
After preprocessing, we have samples that include 46,500 values (93 * 500). This high-dimensional samples, considering the inherent variability of EEG data, cannot be compared directly. Therefore, we need to extract features related to the identity of the subject. There are two main approaches: the first involves classic feature extraction, which relies on predefined formulas derived from expert knowledge in the field. The second approach employs deep learning techniques for feature extraction, allowing models to automatically learn representations from the data.
We select Power Spectral Density (PSD) as the classic feature extraction technique because it effectively transfers signals from the time domain to the frequency domain, capturing unique patterns linked to individual brain dynamics. Consequently, PSD is a widely utilized feature extraction technique in brainwave-based authentication systems (Khalil et al., 2024; Ong et al., 2018; Thomas and Vinod, 2018; Arias-Cabarcos et al., 2023; Chaurasia et al., 2024). PSD applies Fourier transformation to quantify the distribution of signal power across different frequency components (Welch, 1967). The resulting frequency spectrum can then be analyzed either within predefined frequency bands—such as delta, theta, alpha, beta, and gamma—or across the full spectrum, depending on the specific application requirements (Arias-Cabarcos et al., 2023).
For deep learning feature extraction, we used a metric learning approach, which is a type of machine learning paradigm designed to learn a similarity or distance function that maps input data to a feature space where similar instances are closer together and dissimilar instances are farther apart (Kulis et al., 2013). This method is particularly suitable for biometric authentication as it ensures that features of the same identity are clustered closely while features from different identities are well-separated (Schroff et al., 2015; Wang et al., 2017). Metric learning approaches can be categorized into pairwise-based and softmax-based methods. Pairwise-based loss functions optimize embedding distances by focusing directly on pairs or triplets of samples, increasing the number of training instances derived from the dataset. This property makes them particularly effective for small datasets. In contrast, softmax-based methods optimize class separability by mapping embeddings into distinct clusters, typically requiring larger datasets for effective generalization (Horiguchi et al., 2019).
To evaluate the effectiveness of metric learning loss functions in brainwave-based authentication, we used both softmax and pairwise approaches. Although PEERS is significantly larger than all previous brainwave datasets, it is still considerably smaller than face recognition datasets, such as MS-Celeb-1M, which contains over 10 million images from 100,000 identities (Guo et al., 2016). We selected SoftTripleLoss (Qian et al., 2019) and ArcFaceLoss (Deng et al., 2019) as our softmax-based approaches. SoftTripleLoss was chosen for its ability to handle intra-class variations by introducing multiple learnable centers per class, making it particularly effective for data with diverse feature distributions. On the other hand, ArcFaceLoss was selected for its angular margin constraints, which enforce strict inter-class separability and robustness against noise. We selected three pairwise-based loss functions: (1) Triplet Loss (Schroff et al., 2015), which optimizes embeddings by minimizing the distance between an anchor and a positive sample while maximizing the distance from a negative sample, serving as a widely used baseline; (2) LiftedStructureLoss (Oh Song et al., 2016), which leverages all positive and negative pairs within a batch simultaneously, penalizing violations of distance constraints and ensuring stable optimization; and (3) SupConLoss (Supervised Contrastive Loss) (Khosla et al., 2020), a recent and promising method that extends contrastive learning by explicitly using class labels to group all embeddings of the same class closer together while separating embeddings of different classes. This supervision allows SupConLoss to align embeddings more effectively across the entire batch, making it particularly suited for learning robust and discriminative feature representations.
3.4. Feature Comparison
After transferring and reducing brainwave samples from higher to lower dimensions via feature extraction, the next step involves comparing the extracted features and calculating a similarity score to determine verification attempts. Two main approaches are commonly used for this task. The first approach employs distance metrics, and we specifically select cosine similarity and Euclidean distance as they are widely used in biometric studies (Fallahi et al., 2024a; Jiang et al., 2023) and serve as distance functions in pairwise metric learning loss functions. The second approach involves training a binary classifier to distinguish between positive and negative samples for each subject. Positive samples correspond to data from the enrollment session, while negative samples are data from other subjects. However, negative samples used during evaluation must not overlap with those used in training, as this might lead to overestimation of performance by introducing familiarity with the subjects present in the training data—an unrealistic scenario in real-world authentication (Fallahi et al., 2024a, 2023). We select Support Vector Machines (SVM) and Random Forests, as they demonstrate superior performance in prior works (Arias-Cabarcos et al., 2023; Chaurasia et al., 2024), and are widely applied in brainwave authentication studies (Hu et al., 2021; Nagarajan et al., 2023; Shams et al., 2022).
3.5. Threat model
To evaluate the security of brainwave authentication systems as unlocking mechanisms for standalone devices (e.g., PCs, XR devices), we adopt zero-effort attacks (Mansfield and Wayman, 2002) as the primary adversary model, aligning with established biometric authentication standards (National Institute of Standards and Technology, 2023). A zero-effort attack in biometric authentication occurs when an adversary uses their own biometric sample in an attempt to gain unauthorized access to another individual’s account. In this context, the security objective of the biometric system is to minimize the probability that the authentication algorithm incorrectly verifies the adversary’s identity as the legitimate user. Specifically, the likelihood of successful verification of a falsely claimed identity, based on the adversary’s sensor readings, should be negligible. An active impersonation is not feasible at this stage. Because an attacker cannot infer any EEG-related information by observing the victim, nor is it feasible to collect EEG data from the victim without their consent. Even if EEG data were accessed, replay attacks can be prevented by implementing task-dependent brainwave authentication (Lin et al., 2018).
3.6. Performance Metrics
We used the Equal Error Rate (EER) as a summary metric, indicating the point where the False Acceptance Rate (FAR) and the False Rejection Rate (FRR) are equal. The FAR represents the proportion of unauthorized attempts incorrectly accepted by the system, while the FRR refers to the proportion of legitimate attempts incorrectly rejected. EER has been widely used in biometric authentication papers (Rui and Yan, 2018; Ayeswarya and Singh, 2024) to report results and ensure comparability. However, in practical usage, it is preferred to keep FAR at very low rates to ensure system security, even at the cost of a higher FRR. Therefore, we also report the FRR at a FAR threshold of 1% for comparison with other brainwave authentication studies (Arias-Cabarcos et al., 2023; Fallahi et al., 2023; Navaratnam-Blair et al., 2022), as well as at 0.1% [^5] https://pages.nist.gov/800-63-3/sp800-63b.html and 0.01% [^6] https://pages.nist.gov/800-63-4/sp800-63b.html to align with biometric standards.
Table 1. Heatmap of benchmarking EER values for different combinations of feature extraction (PSD, PCA, SupConLoss, ArcFaceLoss) and comparison methods (RF, SVM, Cosine, Euclidean) in brainwave-based authentication, evaluated on 50 test subjects. NTS denotes the number of training and validation subjects used for training deep learning-based feature extractors.
NTS Feature Extractor Output Dimension Comparison Metric
Classifiers Distance Metrics
RF SVM Cosine Euclidean
0 Random Data 93 × 500 49.85 $\pm$ 0.22 49.99 $\pm$ 0.72 49.94 $\pm$ 0.88 49.92 $\pm$ 0.82
Raw EEG 93 × 500 39.49 $\pm$ 9.27 50.07 $\pm$ 0.71 38.07 $\pm$ 10.77 50.45 $\pm$ 13.61
PSD 93 × 129 32.53 $\pm$ 13.97 32.82 $\pm$ 12.23 39.30 $\pm$ 13.28 50.26 $\pm$ 12.28
PSD + Bins 93 × 5 30.54 $\pm$ 12.26 27.30 $\pm$ 14.33 37.52 $\pm$ 13.80 49.55 $\pm$ 12.45
PCA 93 × 5 37.51 $\pm$ 9.72 49.99 $\pm$ 0.86 37.03 $\pm$ 11.46 50.18 $\pm$ 13.15
PSD + PCA 93 × 5 29.16 $\pm$ 14.03 32.03 $\pm$ 12.13 45.47 $\pm$ 13.85 44.14 $\pm$ 15.52
245 Subjects PSD + SoftTripleLoss 128 30.95 $\pm$ 10.61 16.91 $\pm$ 10.52 15.54 $\pm$ 7.81 31.91 $\pm$ 15.07
PSD + ArcFaceLoss 128 32.83 $\pm$ 9.46 17.58 $\pm$ 10.48 18.52 $\pm$ 8.80 36.92 $\pm$ 16.89
PSD + Triplet Loss 128 35.25 $\pm$ 7.98 14.99 $\pm$ 7.57 15.28 $\pm$ 7.82 17.07 $\pm$ 8.38
PSD + LiftedStructureLoss 128 28.11 $\pm$ 11.33 12.67 $\pm$ 7.37 12.19 $\pm$ 7.15 17.34 $\pm$ 9.86
PSD + SupConLoss 128 27.73 $\pm$ 11.20 13.89 $\pm$ 8.36 12.34 $\pm$ 7.83 16.63 $\pm$ 9.58
SoftTripleLoss 128 29.88 $\pm$ 10.41 16.68 $\pm$ 10.29 15.59 $\pm$ 8.48 20.83 $\pm$ 12.18
ArcFaceLoss 128 27.53 $\pm$ 10.93 14.20 $\pm$ 8.34 14.22 $\pm$ 8.20 18.35 $\pm$ 10.96
Triplet Loss 128 35.61 $\pm$ 8.77 13.82 $\pm$ 8.78 13.04 $\pm$ 7.71 14.61 $\pm$ 9.28
LiftedStructureLoss 128 28.97 $\pm$ 11.59 12.41 $\pm$ 8.73 11.38 $\pm$ 7.24 14.25 $\pm$ 11.22
SupConLoss 128 25.62 $\pm$ 12.30 10.91 $\pm$ 6.84 10.75 $\pm$ 7.06 13.38 $\pm$ 9.85
4. Results and Discussion
In order to understand multisession performance, we first require a pipeline that includes a feature extractor and a comparison metric, which we select through rigorous benchmarking. We then investigate important parameters such as test size, time interval, recording hardware, and subject embeddings visualized as distinct clusters through t-SNE plots. Next, to move toward consumer-grade devices, we analyze performance under a reduced number of channels and further contextualize our findings with current state-of-the-art methods and authentication standards to identify potential directions for future research.
4.1. Benchmarking Feature Extraction and Compression
We have introduced the dataset and preprocessing methods; however, selecting suitable feature extraction and compression techniques is essential to finalize the authentication pipeline.
Experiment setup: To perform the benchmark, we need subject-exclusive sets for training and validation of the feature extractor, negative subjects for binary classification feature comparison, and test subjects for evaluation. The deep learning-based feature extraction approach was trained using 245 subjects, with 230 allocated for training and 15 for validation. Meanwhile, 50 subjects were specifically selected as negative samples for the shallow classifier feature comparison. Finally, we randomly selected 50 subjects to serve as the final test set.
Feature Extraction methods: Table 1 presents the results of the benchmark, beginning with the creation of random samples via a uniform distribution matching the size of the original data. This step ensured the absence of biases or misimplementations that could favor lower error results. The random samples achieved an expected baseline EER of approximately 50%. Classic feature extraction methods showed an improvement over this baseline but still yielded relatively high error rates. Among them, PSD features divided into bins demonstrated the best performance with a 27% EER. Transitioning to deep learning-based feature extraction methods resulted in significant performance improvements of 10% points of the EER or more. The best results were achieved by deep learning approaches utilizing raw data instead of preprocessed PSD features. Within these approaches, pair-based methods outperformed softmax-based ones. This advantage was expected due to the relatively larger dataset size compared to previous EEG authentication datasets, which allowed the methods to outperform PSD features. However, the dataset size remains smaller than the extensive datasets used for face detection, which often include thousands of subjects. Consequently, pair-based methods effectively leveraged the available data for better performance.
Comparison Methods: Table 1 demonstrates that SVM and cosine similarity outperform Random Forest and Euclidean distance as comparison functions. However, among the best results, which belong to Supervised Contrastive and Lifted Structure Loss feature extraction methods with raw and PSD input, the cosine similarity function consistently performs slightly better than SVM. The better performance with cosine similarity over SVM arises because deep learning loss functions optimize embeddings using distance metrics like cosine similarity or Euclidean distance in training stage.
Best Performing Authentication Pipeline Selection: The SupConLoss achieved the best results with an error rate of 10.75%, compared to 27.30% for the PSD approach, representing a 16.55% point improvement in the EER. Therefore, we selected SupConLoss as the feature extraction method for its superior performance. For feature comparison functions, cosine similarity, which does not require any prior data, outperformed the binary classification approach that requires 50 subjects as negative samples. Therefore, we selected cosine similarity for feature comparison in the remainder of the paper to apply within further investigations.
Figure 1. Number of sessions available in the test dataset after the enrollment session. Here, $W$ represents a week, $M$ represents a month, and $Y$ represents a year. The value following each symbol indicates the time period it represents, starting at a day after the end of the previous time unit and spanning the duration defined by the symbol ($W$, $M$, $Y$). For example, $W3$ refers to the number of sessions in the third week, beginning at the end of the second week (15th day) and ending on the 21st day. Similarly, $M2$ refers to the number of sessions during the second month, starting at the end of the first month (approximately the 31st day) and covering the next month. Note that $M\approx 30.42$ days and $Y=365.25$ days.
Figure 2. EER vs. Number of Subjects: Each blue dot represents the average of 50 instances of randomly selecting $N$ subjects and calculating the EER. The dot size represents the standard deviation of the EER values across these 50 calculations. The trendline, calculated using ordinary least squares (OLS) regression, minimizes the squared differences between observed data points and the fitted line to reveal any trend between the number of subjects in test set and the average EER.
4.2. Detailed Performance Analysis
With the best-performing pipeline identified, the EER remains significantly lower than the 50% chance level but still reaches 10.75%. This could indicate that brainwaves are not sufficiently unique for authentication or that other underlying issues need to be understood to develop solutions. Therefore, we aim to investigate which parameters influence multi-session authentication more deeply by addressing the following questions:
  1. (1)
    How does altering the number of participants in an EEG-based authentication test set affect the overall comparability of the evaluation? To investigate this, we analyze how changing the number of test subjects impacts the EER in different scenarios.
  2. (2)
    How does EEG-based authentication performance change over time, and what does this suggest about brainwave pattern stability? To explore this, we examine how EER is influenced by varying verification and enrollment intervals.
  3. (3)
    How effectively are subjects clustered based on their embedding space derived from the feature extractor? This explores the visualization of embedding space.
  4. (4)
    How does different recording hardware affect the performance of brainwave-based authentication, and what are the best practices for ensuring reliability? This question is investigated by measuring the EER on hardware-specific and generalized models as well as evaluating cross-hardware authentication scenario.
Increasing the amount of test data provides deeper insights and allows for a more comprehensive evaluation of the multi-session authentication system. In the previous step, we selected cosine similarity as the feature comparison function. Consequently, the 50 subjects previously designated as negative samples for binary classification can now be incorporated into our evaluation, as they were not utilized in any part of the current pipeline. This inclusion allows us to expand the test set to 100 subjects, enabling a more robust assessment of system performance. We observe that using SupConLoss and cosine similarity results in an EER of 10.78% ± 6.67% on the 100 test subjects, showing only a marginal 0.03% increase relative to the 50-subject test set. We could also use the data from these 50 subjects to improve the feature extractor, which would likely result in a slightly lower EER. However, this approach would limit our ability to investigate scenarios with fewer sessions across different time intervals or a smaller number of test subjects per headset.
Figure 3. Relationship between the time interval between sessions and EER. $D$ represents a day, $W$ a week, $M$ a month, and $Y$ a year. The value after each symbol indicates the start time interval (except for days). For example, $W1$ represents sessions with a time interval of 8–14 days after the first session.
4.2.1. Effect of the Number of Test Subjects on EER:
When comparing the results of different experiments using metrics such as EER, number of subjects in the test is typically reported as metadata. A larger test set generates more impostor scores in zero-effort scenarios, making it valuable to investigate the relationship between the number of subjects in the test set and the resulting error. This analysis provides insights for better comparisons of results across different studies or our follow-up experiments with smaller test sets, such as when evaluating performance using headset-specific recording. To achieve this, we randomly select $N$ subjects, where $N$ ranges from 2 to 99 (starting at $N=2$ since at least one genuine user and one impostor are required). For each value of $N$, we compute the EER, repeat the procedure 50 times, and report the average EER per $N$.
Figure 2 presents the results of the experiment. The analysis revealed no significant relationship between the number of subjects and the EER. These findings align with observations in face recognition systems. For example, Friedman et al. (Friedman et al., 2022) investigated how test size affects identification and authentication tasks. They reported a linear decline in identification performance with the logarithm of enrollment size but found that the EER for authentication tasks remained stable. This stability occurs because the EER is based on the Receiver Operating Characteristic (ROC) curve, which balances the FAR and FRR derived from impostor and genuine score distributions. While larger subject pools better define these distributions, they do not significantly influence the ROC. Therefore, we can conclude that having a higher number of subjects in the test set does not necessarily make authentication more challenging. However, both scenarios indicate that test sets with a small number of participants exhibit high standard deviations in EER, which can lead to unreliable or misleading results. This underscores the importance of using larger test sets to obtain more dependable evaluations.
Figure 4. t-SNE Visualization of user identity embeddings, where a color is assigned to each of the 100 subjects in the test set. Due to the large number of subjects, some colors may repeat.
4.2.2. Correlation Between EER and Time Interval:
To investigate the permanence, we first visualize the distribution of test session intervals in Figure 3. The test data includes a total of 1,699 unique sessions available for 100 test subjects, where the first session (or the oldest session) per subject is considered as the enrollment session (100 sessions). The results show that the highest concentration of sessions occurred within the first week, with 129 sessions (8%) recorded. By the end of the first month (weeks 0 to 3), this number increased to 560 sessions (35%), indicating that one-third of all sessions were conducted in the first month. The second month (weeks 4 to 7) accounted for 398 sessions (24.8%), while the third month (weeks 8 to 11) added 201 sessions (12.5%). Beyond the third month, session frequency declined further, with 182 sessions (11.3%) recorded from Month 4 to the end of the first year. In the second year, session numbers were 89 sessions (5.5%), and remained similar at 71 sessions (4.4%) in the third year. The remaining years combined accounted for 4.8%.
To investigate the correlation between time intervals between sessions and authentication error, we extracted all possible pairs of sessions within specific time intervals and calculated the EER for each. Figure 3 shows the results, demonstrating a positive correlation between time interval and error rate, as expected. Verifying identity becomes harder over time. The EER after one day is 7.7%, increasing to 8.6% after seven days. However, within the second month, the EER rises to 10.4%, which remains lower than the previously reported average of 10.78%, indicating stable performance over a reasonable time interval. Beyond the second month, the EER gradually increases, reaching 19.69% after one year. Performance remains relatively stable in the second and third years. However, in the fourth and fifth years, the performance deteriorates significantly, approaching random baseline levels with EER values of 40% and 75%, respectively. Therefore, it is necessary to update the enrollment set at shorter time intervals. This can be achieved by using verification data from successful login attempts to update the enrollment set, ensuring that the process does not impose additional overhead on the user.
Figure 5. t-SNE visualization of embeddings, where each color represents a different EEG recording headset used for data collection.
4.2.3. Visualize Embedding Space:
To better understand how well users are represented in distinct clusters, we transform the 128-dimensional embeddings to a 2D space using t-Distributed Stochastic Neighbor Embedding (t-SNE) and visualize the results. While some noise is present, the Figure 4 shows that the metric learning approach successfully separates subjects into distinguishable clusters, demonstrating the model’s ability to capture individual differences. However, we observe two separate colonies within the embedding space, leading us to hypothesize that these colonies could be due to differences in the hardware used during data capture. This hypothesis is confirmed by Figure 5, which reveals that the data captured using BioSemi hardware is located separately from the rest, further illustrating the significant influence of hardware on the embedding distribution. The effect of hardware on the results will be discussed in the next section.
Figure 6. t-SNE visualization of embeddings for 20 randomly selected subjects, with each subject assigned a unique color for a clear distinction. Enrollment samples are depicted as circles without borders, while verification samples are represented by various shapes with black borders. Also, each unique shape corresponds to a distinct session for verification samples.
Figure 7. Magnified view of the right-middle region of Figure 6, highlighting the green subject. Each session for the highlighted subject is assigned a unique color, while all other subjects are shown in gray for context.
To investigate the embedding space while accounting for color and spatial limitations, we randomly selected 20 subjects to prevent overcrowding the embedding space with excessive data. [^7] BioSemi subjects were excluded to avoid large gaps in the plot and enhance detail visibility. Figure 6 illustrates the embedding space for these selected subjects, demonstrating that the feature extraction method effectively identifies subject identities. However, a session effect is evident, where samples from the same session tend to form distinct clusters. To better visualize this phenomenon, the right-middle region of Figure 6 is magnified in Figure 7, with separate colors indicating different sessions, further confirming the session-based clustering. The session effect on clustering could primarily arise from environmental noise, changes in hair length, slight variations in the positioning of brainwave channels on the head, hardware differences, and natural changes in brainwave patterns over time due to factors such as aging, personal experiences, or mental states (e.g., stress or focus levels) throughout sessions.
4.2.4. Impact of Different Headsets on Performance
Due to differences in recording hardware technologies and channel positions in EEG recording devices, further investigation is required to understand their effect on brainwave authentication. Specifically, we aim to address three questions: (1) To what extent can a model trained on one hardware be used for the authentication of user data recorded by another hardware? (2) Can models trained on multiple hardware types outperform headset-specific models? (3) What is the performance of cross-device authentication when enrollment and verification data come from different hardware?
Cross Model: To address the first question, we trained models using data from each specific headset and grouped the test data based on hardware to evaluate performance across models. The results, summarized in Table 2, show that headset-specific models perform well only when tested on data from the same device. For example, the model trained on HydroCe achieves an EER of 12% on HydroCe test data but 27% on Geodisi test data. To further investigate why some hardware-specific models perform better on their own data, we summarized the training and testing data in Table 4, which reveals that the number of unique subjects in the training set has a stronger correlation with the EER. Therefore, weaker performance in some models could be related to the lower number of subjects available for learning features.
General Model: We trained a general model using data from all devices. The results, presented in Table 2, show that the general model slightly outperforms hardware-specific models even on their own data. However, a follow-up question arises: does training a model on independent hardware-specific datasets contribute to better general models, or does the general model benefit from learning cross-device patterns in the current training dataset? To explore this, we created a smaller train dataset by retaining only the most frequent hardware per subject, ensuring that each subject exclusively has data recorded by a single device. The results in Table 3 indicate that the general model outperforms the Geodisi and BioSemi hardware-specific models but performs slightly worse than the HydroCe-specific model. Therefore, to some degree, we can conclude that combining smaller datasets with electrodes in the same positions can be beneficial for brainwave authentication, especially when large datasets are not available.
Cross Hardware: In real-world scenarios, individuals using brainwave-based authentication systems should be able to rely on various hardware devices. Therefore, subjects should still be authenticated when using different recording devices. Two key observations provide insight into this problem. First, the general model’s performance, when trained on a hardware-independent dataset, is presented in Table 3. It shows an error rate more than twice as high compared to Table 2, where the model was trained on a dataset including subjects with recordings from different hardwares. This suggests that providing cross-hardware samples during feature extractor training is essential for improving performance. Second, we filtered genuine scores based on unique enrollment and verification hardware pair combinations in the test dataset and calculated the EER for cross-hardware scenarios. The results show an EER of 13.25% for Geodisi-HydroCe pairs and 48.2% for BioSemi-HydroCe pairs and there were no HydroCe-BioSemi pairs in the test set. These results can be explained by the fact that only 7 subjects in the training set had cross BioSemi-HydroCe sessions, compared to 125 subjects who had cross Geodesic-HydroCe sessions. This could also explain the formation of separate clusters in the t-SNE visualization shown in Figure 5, where BioSemi forms a distinct cluster.
Table 2. Table showing EER (%) with standard deviation ($\pm$ Std). The first column (D) lists the datasets, while the first row (M) lists the models. For example, the entry in the second row and third column (27.48 $\pm$ 10.39) represents the EER for the HydroCe model (train based on HydroCe data) evaluated on the Geodisi dataset. ”All” indicates results obtained using all data or models trained on data from all headsets.
D/M Geodisi HydroCe BioSemi All
Geodisi $14.50\pm\text{\scriptsize 8.33}$ $27.48\pm\text{\scriptsize 10.39}$ $37.81\pm\text{\scriptsize 10.87}$ $13.37\pm\text{\scriptsize 7.13}$
HydroCe $30.34\pm\text{\scriptsize 11.15}$ $12.15\pm\text{\scriptsize 7.86}$ $34.63\pm\text{\scriptsize 10.37}$ $12.26\pm\text{\scriptsize 9.12}$
BioSemi $42.11\pm\text{\scriptsize 7.10}$ $39.74\pm\text{\scriptsize 9.22}$ $23.28\pm\text{\scriptsize 8.60}$ $20.10\pm\text{\scriptsize 9.62}$
All $27.93\pm\text{\scriptsize 12.16}$ $23.45\pm\text{\scriptsize 13.50}$ $30.71\pm\text{\scriptsize 12.42}$ $10.78\pm\text{\scriptsize 6.67}$
Table 3. Table showing EER (%) with standard deviation ($\pm$ Std). The first column (D) lists the datasets, and the first row (M) lists the models, where the training data for each subject is limited to their most frequently used hardware.
D/M Geodisi HydroCe BioSemi All
Geodisi $21.93\pm\text{\scriptsize 10.40}$ $28.85\pm\text{\scriptsize 11.79}$ $34.43\pm\text{\scriptsize 7.46}$ $17.82\pm\text{\scriptsize 9.32}$
HydroCe $35.08\pm\text{\scriptsize 9.63}$ $12.49\pm\text{\scriptsize 9.23}$ $33.92\pm\text{\scriptsize 10.12}$ $13.82\pm\text{\scriptsize 9.60}$
BioSemi $43.99\pm\text{\scriptsize 5.94}$ $40.57\pm\text{\scriptsize 6.40}$ $25.52\pm\text{\scriptsize 8.15}$ $22.03\pm\text{\scriptsize 8.98}$
All $32.56\pm\text{\scriptsize 10.54}$ $26.42\pm\text{\scriptsize 16.29}$ $29.23\pm\text{\scriptsize 10.41}$ $24.37\pm\text{\scriptsize 17.22}$
Table 4. Summary of train datasets by hardware type. The table lists the number of sessions and subjects for each hardware type, with the corresponding percentages shown in parentheses.
Hardware Sessions Subjects
Train Set: Full Dataset
   Hardware HydroCe 2038 169
   Hardware Geodisi 1016 127
   Hardware BioSemi 945 66
Train Set: Filtered Dataset
   Hardware HydroCe 1751 131
   Hardware Geodisi 435 38
   Hardware BioSemi 905 61
Table 5. Tradeoff between the number of samples (SN) and EER. Increasing the number of samples reduces the EER but increases the verification time, as shown in the table.
SN 1 2 4 8 16 32
EER 10.78 9.04 7.98 7.38 7.04 6.87
Time 1s 2s 4s 8s 16s 32s
Table 6. Tradeoff between the number of enrollment sessions (ES) and EER when using 1 verification sample (1V) and 4 verification samples (4V). The table also includes the number of available subjects (NAS) used in the evaluation.
ES 1 2 3 4 5 6
1V 7.67 6.01 5.80 6.05 5.38 6.70
4V 4.23 2.67 2.78 2.67 2.23 2.90
NAS 100 99 99 98 91 75
Table 7. EER (%) when using the same channels as consumer-grade devices. $E_{x}$ denotes the number of enrollment sessions, and $V_{x}$ denotes the number of verification sessions.
ES Baseline 4V E1 V1 E1 V4 E2 V1 E2 V4
Emotiv 13.97 10.74 10.44 6.01 8.79 4.48
DSI-VR300 16.55 12.70 13.49 8.56 11.94 6.87
Muse 2 21.94 17.74 18.07 12.42 17.58 11.19
4.3. Enhancing Performance
Samples from the same session are relatively close in the embedding space, but in some cases, they are more dispersed. For example, the pink triangle and green plus in Figure 6. Therefore utilizing multiple verification samples can improve results, albeit at the cost of additional verification time. Table 5 illustrates the tradeoff between the number of verification samples, EER, and verification time. The results show that using more samples reduces the error, but improvements beyond four samples show a slower pace in reducing error. Considering the time cost of approximately one second per sample, using four samples appears to be a reasonable balance for the averaging strategy.
Moreover, while comparing the enrollment session to all verification sessions provides valuable insights into brainwave authentication, real-world scenarios typically involve the first verification occurring immediately after enrollment. Additionally, verification sessions can potentially be used as supplementary data to enhance the enrollment process in the next verification attempt. Therefore, in Table 6, we consider the first session as enrollment and the second as verification and report the results. We observe an EER of 7.67% based on single verification sample and 4.23% for four verification samples, showing significant improvement compared to the previous 10.78% and 7.98%. Additionally, we investigate the effect of multiple enrollment sessions. The results suggest that having two enrollment sessions improves EER to 6.01% and 2.67%, but further enrollment sessions show diminishing returns. It is worth mentioning that the average time interval between two enrollment sessions was 28 days. Among the subjects, 88 had a time interval of 10 days or less, while 32 had session intervals of only one or two days.
4.4. Practicality with Consumer-Grade Channels
Despite advancements in brainwave sensors in recent years, consumer-grade devices still have a limited number of sensors. Therefore, we selected three devices to evaluate performance by focusing exclusively on the electrodes specific to each device. The selected devices include the Emotiv Epoc X, which features 14888AF3, F7, F3, FC5, T7, P7, O1, O2, P8, T8, FC6, F4, F8, AF4 electrodes and is priced at 999 dollars, and the Muse 2, equipped with 4999TP9, AF7, AF8, TP10 electrodes and available at 269.99 euros, offer a budget-friendly options for consumer use. Additionally, we considered the DSI-VR300101010FCz, Pz, P3, P4, PO7, PO8, Oz, which includes 7 electrodes and costs approximately 10,000 dollars. This device, though expensive, has the potential to become more affordable in the future and could be integrated into XR devices.
We selected the same electrodes as used in the devices and trained a model for each device. Table 7 represents the results with a lower number of channels. The baseline EER increases from 10.78% to 13.97%, 16.55%, and 21.94% for Emotiv Epoc X, DSI-VR300, and Muse 2, respectively. However, considering the 93 channels in our main dataset and reducing them to 14, 7, and 4 channels indicates some redundant information in the original channels due to EEG correlations across electrode recordings. Moreover, using averaging techniques and multiple enrollment sessions can further improve the results.
Table 8. Summary of state-of-the-art brain biometric verification solutions:
Publication Subjects Sess. Sess. Interval Verification time public dataset open source EER (%)
Das et al.  (Das et al., 2016), 2016 50 150 up to 49 days 30s $\times$ $\times$ 10
Wu et al.  (Wu et al., 2018), 2018 40 80 30 days 6s $\times$ $\times$ $5.75<\text{EER}<7.07$
Maiorana et al.  (Maiorana and Campisi, 2017), 2018 45 270 up to 36 months 5s $\times$ $\times$ $6.6\leq\text{EER}\leq 10.7$
Maiorana (Maiorana, 2021), 2021 45 225 up to 15 months 5s $\times$ $\times$ $8\leq\text{EER}\leq 10.7$
Debie et al. (Debie et al., 2021), 2021 54 108 different days [^3] The paper did not specify the exact number of days, so we used the same terminology. However, it typically refers to a very short time interval. $1s-4s$$\times$ $1.32<\text{EER}<2.22$
Chaurasia et al. (Chaurasia et al., 2024), 2024 54 108 different days [^3] The paper did not specify the exact number of days, so we used the same terminology. However, it typically refers to a very short time interval. 1s 11.99
Proposed Method 345 6013 up to 66 months 1s 10.78
Proposed Method 345 6013 up to 66 months 4s 7.98
4.5. Comparative Analysis
We first evaluate the proposed method against state-of-the-art brainwave authentication approaches. Subsequently, we analyze its performance relative to international biometric standards to provide context and clarify the significance of the results.
4.5.1. Compare with Brainwave Authentications Systems
Table 8 summarizes the state-of-the-art in multi-session brainwave authentication studies. As shown, previous studies typically involve 40–54 subjects and 80–270 sessions, which is significantly lower than our dataset of 345 subjects and 6013 sessions. From the perspective of time intervals between sessions, prior works vary: some datasets span a few days, such as those in (Chaurasia et al., 2024; Debie et al., 2021), which report inconsistent EER values ranging from 1.32% to 11.99%. Others span 1–2 months with EER values between 5.75% and 10%(Das et al., 2016; Wu et al., 2018), and the study by Maiorana et al.(Maiorana, 2021; Maiorana and Campisi, 2017) includes intervals of up to 36 months with EER values between 6.6% and 10.7%. However, our dataset includes time intervals of up to 66 months with EER values of 7.98% and 10.78%.
While we have already demonstrated that a lower number of subjects or sessions does not necessarily make authentication evaluation easier, we observe that evaluation with a lower number of subjects could have higher variability, making it more susceptible to overfitting. We also observed that longer time intervals between sessions can significantly increase the error rate. Therefore, in terms of results, we can conclude that our findings are comparable to or outperform the state-of-the-art, despite using a more challenging dataset and shorter verification times. However, only Debie et al. (Debie et al., 2021) show a lower EER compared to our result, even when comparing our result after a day (7.7%), where they train a deep model network per subject based on a single session. While this approach is not scalable as it requires training a CNN per subject, it is not obvious how the model could learn session variants while learning only from a single session and evaluated by the second one, and the source code was not public. But Chaurasia et al.(Chaurasia et al., 2024) used the same dataset and reported 20.82% EER based on triplet loss function and Euclidean distance, and 11.99% based on PSD as feature extraction and SVM as feature compression, which shows the data was not enough to take advantage of deep learning even based on multiple subjects. Additionally, our study is the only one to use a publicly available dataset with time intervals exceeding a few days. It is also the only one, except Chaurasia et al.(Chaurasia et al., 2024), to release its source code on multi-session brainwave authentication.
Table 9. Performance metrics for different scenarios with detailed enrollment and verification session information.
Scenario Metrics (%)
Enrollment Sess. Verification Sess. Verification Samples EER FMR 1% FMR 0.1% FMR 0.01%
First All Others 1 $10.78\pm 6.67$ $52.82\pm 24.54$ $79.77\pm 18.41$ $91.97\pm 11.43$
First All Others 4 $7.98\pm 7.00$ $40.97\pm 29.46$ $67.56\pm 26.29$ $81.86\pm 21.59$
First Second 1 $7.67\pm 7.28$ $45.77\pm 33.58$ $73.51\pm 29.72$ $84.76\pm 23.37$
First Second 4 $4.23\pm 6.39$ $35.10\pm 38.63$ $56.06\pm 41.13$ $66.15\pm 38.80$
First Two Third 1 $6.01\pm 5.29$ $35.98\pm 28.36$ $66.49\pm 27.72$ $81.57\pm 22.49$
First Two Third 4 $2.67\pm 3.87$ $22.67\pm 33.28$ $44.78\pm 38.92$ $57.59\pm 38.31$
First Two Third 16 $1.41\pm 2.70$ $21.13\pm 37.59$ $31.87\pm 41.42$ $38.96\pm 42.80$
4.5.2. Comparison with Biometric Industry Standards
Brainwave authentication is still in its early stages, making it crucial to assess its current progress and identify the gaps for real-world implementation. For a biometric system to be deployed in practical applications, it must adhere to international industrial biometric standards. These standards emphasize a very low FAR to ensure security, while maintaining a reasonable FRR to keep the system usable for legitimate users. According to guidelines from NIST (2023)/ISO [^11] https://pages.nist.gov/800-63-3/sp800-63b.htmland the European Border Guard Agency Frontex [^12] https://www.frontex.europa.eu/assets/Publications/Research/Best_Practice_Technical_Guidelines_ABC.pdf, biometric systems are required to operate at FAR $\leq 0.1\%$ (1 in 1,000). Stricter recommendations from FIDO [^13] https://fidoalliance.org/specs/biometric/requirements/Biometrics-Requirements-v4.0.1-fd-20240522.pdf and the upcoming NIST (August 2024)/ISO 2024 standards [^14] https://pages.nist.gov/800-63-4/sp800-63b.html suggest a FAR $\leq 0.01\%$ (1 in 10,000). For usability, FRR standards propose a maximum FRR of $\leq 5\%$, allowing at least 19 successful logins out of 20 attempts for genuine users.
Table 9 presents the FRR at low False FAR of 1%, 0.1%, and 0.01%. The results show higher FRR. For instance, with two enrollment sessions and a verification time of 4 seconds, the FRR is 44.78% at FAR 0.1% and 57.59% at FAR 0.01%, meaning approximately 11 and 8 successful logins out of 20 attempts for legitimate users, respectively. However, we also observed very high standard deviations, indicating that the system works well for some subjects but performs worse for others. For example, under the same scenario, the FRR is less than or equal to 20% for 41 and 26 subjects at FAR 0.1% and is less than or equal to 5% for 30 and 14 subjects at FAR 0.01%. These observations highlight the need for further investigation at the subject level.
While our experiments show that brainwave authentication does not yet meet international standards, our previous experiment on the impact of headsets on performance (Sec.4.2.4) revealed that the headset with a higher number of subjects yielded a lower error rate, and we also know that biometrics that meet these standards typically employ datasets with thousands of subjects(Guo et al., 2016). Therefore, we conducted an experiment to explore the relationship between the number of subjects used for training the feature extractor and the error rate. The results, presented in Figure 8, indicate a linear relationship between the logarithm of the number of subjects and the EER. Based on this relationship, we predict that if 1,024–2,048 subjects were available for training, the EER could drop below 0.5%, which would satisfy the required standards (Fallahi et al., 2024a).
Figure 8. Relationship between the number of subjects used to train the feature extractor and the EER. The plot shows experimental results (blue points with error bars) and a linear prediction (dashed line) based on these results.
5. Limitations
Our study on the effective parameters influencing EEG-based authentication performance in a multi-session scenario is subject to certain limitations, discussed below. First, our stimuli were limited to a general text-based memorization task. Ideally, multiple task types should have been included to allow for performance comparisons and identification of the best-performing task. Second, we simulated consumer-grade devices by selecting the same number of channels in the same locations as those found in such devices. However, we did not have a measure to directly compare the quality of sensors. Additionally, we did not consider downsampling the data to simulate sensor quality, as downsampling served as a noise reduction method in multi-session authentication (Maiorana, 2021). It is worth noting that all other multi-session studies in the literature (Maiorana and Campisi, 2017; Debie et al., 2021; Wu et al., 2018; Chaurasia et al., 2024) also employ medical-grade data. Finally, we had access to only a single dataset that included a notable number of sessions and subjects. However, best practices in state-of-the-art face recognition research recommend training on multiple large datasets and testing on an independent dataset to ensure a more robust evaluation (Kim et al., 2022). Achieving this requires massive amounts of data to extract distinctive features while accounting for subject, environmental, and hardware-specific effects. For example, Kim et al. (Kim et al., 2022) used three datasets comprising over 350,000 subjects and more than 14.9 million images to train a generalized model capable of evaluation on a few independent datasets.
6. Conclusion and Future Work
This study aimed to investigate the factors influencing the performance of brainwave authentication over multiple sessions. We conducted evaluations using the largest multi-session brainwave authentication dataset available to date, which includes 345 subjects and 6,013 sessions. Our findings highlight that a pairwise deep learning approach for feature extraction combined with cosine similarity for feature comparison is the most effective method for processing brainwave data. Interestingly, we observed that larger test sets do not necessarily make authentication more difficult; instead, they reduce the uncertainty of the results. Authentication performance remained relatively stable during the first week, with error rates increasing only slightly in the second month. However, error rates could increase by up to 2.5 times after one year, suggesting that the enrollment dataset should be periodically updated to maintain performance. We also examined the impact of recording hardware on results, emphasizing the importance of including data from different headsets in training the feature extractor. Additionally, we investigated the trade-offs affecting performance and showed that increasing the number of enrollment sessions and verification time can help reduce the EER, though only up to a certain point and at the cost of time. Our findings also demonstrate that brainwave authentication remains feasible even with a reduced number of EEG sensors. Finally, we demonstrate that our results are competitive with the state-of-the-art in brainwave authentication, investigated with stronger assumptions in simpler scenarios. However, achieving industrial standard-level performance would require collecting data from a larger number of subjects.
For future work, there is an urgent need for a public multi-session dataset with more than 1,024 subjects to advance research in brainwave authentication. This dataset should be collected using consumer-grade devices in multiple sessions, and incorporating multiple devices in the same session would be valuable for studying the cross-headset authentication challenge. On the methodological side, further research should focus on developing deep learning architectures that better capture the unique features of brainwave data, as well as effective signal preprocessing approaches to remove noise. Adopting an open-source approach would enable researchers to build on each other’s work, promoting reproducibility and collaboration rather than isolated, non-reproducible studies.
Acknowledgements.
This work was funded by the Topic Engineering Secure Systems of the Helmholtz Association (HGF) and supported by KASTEL Security Research Labs, Karlsruhe, and Germany’s Excellence Strategy (EXC 2050/1 ‘CeTI’; ID 390696704). Portions of the research in this paper use the FERET database of facial images collected under the FERET program.